OpenStudy (anonymous):

i forgot how to simplify and cant get 14/8 * 17/21

jimthompson5910 (jim_thompson5910):

14/8 * 17/21 = (14*17)/(8*21) = 238/168 = 17/12 So 14/8 * 17/21 = 17/12

OpenStudy (anonymous):

238/168 = 119/84

OpenStudy (anonymous):

how tdo u simplify 119/84

OpenStudy (anonymous):

you can't..

OpenStudy (anonymous):

other answer received was 238/68 = 17/12

OpenStudy (anonymous):

that one is right.

OpenStudy (anonymous):

how

jimthompson5910 (jim_thompson5910):

@shaneryco you can divide each term by 7

OpenStudy (anonymous):

its 238/168, then divide top and bottom by 14.

OpenStudy (anonymous):

what bout negatives -12/20 *42/72

OpenStudy (anonymous):

sorry 12/30

OpenStudy (anonymous):

i mean -12/30*-42/72 how do you multiply the negative fractions

jimthompson5910 (jim_thompson5910):

-12/30*-42/72 = (-12*-42)/(30*72) = 504/2160 = 7/30 So -12/30*-42/72 = 7/30

OpenStudy (anonymous):

so to find simplest form where do you start. I mean 9 and go backwards. in other words how did u arrive at 7 for numerator

jimthompson5910 (jim_thompson5910):

I divided both numerator and denominator of 504/2160 by 72 (which is the GCF of the two numbers) to get 7/30

OpenStudy (anonymous):

http://www.difranco.net/cop2220/op-prec.htm